OpenStudy (anonymous):

Find the vertex, focus, and directrix of the parabola and sketch its graph. Use a graphing utility to verify your graph. y2 − 8y − 4x = −12

OpenStudy (anonymous):

You don't have to graph it for me just explain to me how to get the vertex because i keep getting the wrong answer :-(

OpenStudy (freckles):

do you know how to complete the square ?

OpenStudy (freckles):

like y^2-8y+? what would the question mark need to be to make y^2-8y+? a square

OpenStudy (anonymous):

16

OpenStudy (freckles):

right \[y^2+ky+(\frac{k}{2})^2=(y+\frac{k}{2})^2 \\ y^2-8y+(\frac{8}{2})^2=....\] so y^2-8y+16=?

OpenStudy (anonymous):

4x-12

OpenStudy (freckles):

...we aren't looking at your equation right now

OpenStudy (freckles):

I'm asking you to write y^2-8y+16 as a square right now

OpenStudy (anonymous):

(y-4)^2

OpenStudy (freckles):

ok cool stuff let's look at your equation now so we need to add 16 to both sides of our equation

OpenStudy (freckles):

so that we can write that (y-4)^2 somewhere on the left hand side

OpenStudy (freckles):

\[y^2-8y-4x=-12 \\ y^2-8y+16-4x=-12+16 \\ \\ \text{ and it looks like you were wanting to go ahead and add } 4x \\ \text{ \to both sides } \\ \text{ \let's do that } \\ y^2-8y+16=4x-12+16 \\ (y-4)^2=4x-12+16\]

OpenStudy (freckles):

so anyways you can simplify -12+16

OpenStudy (freckles):

and factor whatever coefficient of x is out of the terms on the right hand side

OpenStudy (anonymous):

(y-4)^2=4(x+1) is the equation then?

OpenStudy (freckles):

right

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

so you are able to spot the vertex?

OpenStudy (anonymous):

(-1, 4)

OpenStudy (freckles):

ok cool that's it

OpenStudy (anonymous):

i didn't know we had to add 16 to both sides so that's where i was going wrong oops

OpenStudy (freckles):

yeah in order for equation to remain the same equation you must do to one side as you would do to the other

OpenStudy (freckles):

like for example you know x=x is true but if you add 16 to one side then the equation is false x+16=x is not true

OpenStudy (anonymous):

that makes sense now

OpenStudy (freckles):

but wee know that if x=x is true then x+16=x+16 is true :)

OpenStudy (freckles):

x+16=x+16 is true anyways this was just an example

OpenStudy (freckles):

here is another if x=y is true, then x+16=y+16 is true

OpenStudy (freckles):

anyways good luck